OpenStudy (anonymous):

Factor out the greatest common factor (GCF) 15^{2n}+9^{n+1}-3^{n}

OpenStudy (anonymous):

Hope this helps, sorry if it doesn't.. http://www.wolframalpha.com/input/?i=15^%282n%29+%2B+9^%28n%2B1%29+-3^n

OpenStudy (anonymous):

you have to decompose 15, 9 and 3 as a product of prime factors and then assemble them for example \[15^{2n}=3^{2n}*5^{2n}\]

OpenStudy (anonymous):

\[3^{2n}*5^{2n}+3^{n+1}*3^{n+1}-3^{n}\] but \[3^{n+1}=3*3^{n}\] and \[3^{2n}=3^{2}*3^{n}\] so u can put \[3^{n}\] as (GCF) the solution is \[3^{n}*(3^{2}*5^{2n}+3*3^{n+1}-1)\]